# Measurement of the mean excitation energy of liquid argon

### v2

@@ 5.1 Scintillator and camera
To explain the discrepancy between our data and the expected shape of the Bragg peak, we suggest that the scintillation spectrum of chromox shifts with the dE/dx of protons. Such a shift has been observed when comparing emission from x-ray and alpha irradiation in zinc oxide and zinc sulfide scintillators [13], in which the highly ionizing alphas result in a spectrum shifted roughly 10 nm down. Since we observed the chromox emission through both an infrared filter (“hot mirror”) and a red filter in our camera, we are sensitive to the convolution of the emission spectrum and transmission efficiency. A similar 10 nm shift would cause a 50% change in the signal (see figure 7) because of the sharp edge of the infrared filter, while a 20 nm shift would cause a 200% change.
@@
Since we lack specific data for either the spectral shift or the scintillator decay time as a function of dE/dx, we make a generic assumption that there is one efficiency for low-dE/dx particles and another for high-dE/dx particles, with some transition region between these. We allow the fit to determine the cutoff dE/dx, the width of the transition, and the ratio of efficiencies. We limit the ratio of efficiencies to between 1:1 and 1:3, corresponding to a spectral shift in the chromox of between zero and 20 nm. The data has relatively little power to simultaneously constrain all three parameters, leading to a large systematic uncertainty. The fit prefers an abrupt transition at 60 MeV/cm with an efficiency ratio of 1:2.3, but given the degeneracies between parameters, these specifics should not be taken too literally.
@@
Changes in the scintillator attenuation length are almost exactly degenerate with an overall left-right shift of the predicted light curve, as it is equivalent to adding or subtracting material in the protons’ path. We consider the extreme cases of all of the light having an attenuation length of 0.05 cm and all of it having an attenuation length of 0.125 cm. The difference in inferred I-value between these extremes is 0.7 eV. We take the midpoint as the nominal result, assume a uniform PDF over the range and claim an uncertainty of 0.2 eV.
@@ 5.4 Multiple Coulomb scattering
which is rapid to compute at the expense of accuracy. It predicts a longer projected range which gives an I-value 5.5 eV lower than inferred from the EMZ option. We do not use this rather different result except to note that experiments which rely on absolute particle ranges should certainly use EMZ and not the default Opt0.
@@ 6 Conclusions
Using our recommended value rather than the old values of 188 eV or 187 eV shifts energy reconstruction in liquid argon TPCs used for GeV-scale neutrinos by around 0.4%, with the exact shift depending on the analysis method. For instance, the dE/dx of a minimum-ionizing particle is reduced from 1.508 MeV g-1 cm-2 to $(1.501\pm 0.002)$ MeV g-1 cm-2. The inferred value of $\Delta m^{2}_{32}$ is directly proportional to reconstructed energy. As DUNE projects a sensitivity to this parameter of around 0.5% (depending on exposure) [32], is it important that a reliable value and uncertainty for the mean excitation energy of liquid argon be used.

### v4

@@ 5.1 Scintillator and camera
To explain the discrepancy between our data and the expected shape of the Bragg peak, we suggest that the scintillation spectrum of chromox shifts with the dE/dx of protons. Such a shift has been observed when comparing emission from x-ray and alpha irradiation in zinc oxide and zinc sulfide scintillators [13], in which the highly ionizing alphas result in a spectrum shifted roughly 10 nm down. Since we observed the chromox emission through both an infrared filter (known as a “hot mirror” by photographers) and a red filter, both part of the camera, we are sensitive to the integral of the product of the emission spectrum and transmission efficiency. A similar 10 nm shift would cause a 50% change in the signal (see figure 7) because of the sharp edge of the infrared filter, while a 20 nm shift would cause a 200% change.
@@
Since we lack specific data for either the spectral shift or the scintillator decay time as a function of dE/dx, we make a generic assumption that there is one efficiency for low-dE/dx particles and another for high-dE/dx particles, with some transition region between these. We allow the fit to determine the cutoff dE/dx, the width of the transition, and the ratio of efficiencies. We limit the ratio of efficiencies, low-dE/dx:high-dE/dx, to be between 1:1 and 1:3, corresponding to a spectral shift in the chromox of between zero and 20 nm. The data has relatively little power to simultaneously constrain all three parameters, leading to a large systematic uncertainty. The fit prefers an abrupt transition at 60 MeV/cm with an efficiency ratio of 1:2.3, but given the degeneracies between parameters, these specifics should not be taken too literally.
@@
Changes in the scintillator attenuation length are almost exactly degenerate with an overall left-right shift of the predicted light curve, as it is equivalent to adding or subtracting material in the protons’ path. For example, for an arbitrarily short attenuation length, any proton that reaches the front face produces a visible signal; protons that continue farther produce no additional signal. At the other extreme, for an arbitrarily long attenuation length, light production from all depths contributes, meaning protons with enough energy to reach the back face produce the greatest signal. This situation is nearly equivalent to having a half-thickness of scintillator of extra material in the protons’ path. (The attenuation length also changes the overall normalization of the signal, but this is a free parameter in our fit, and so after fitting, only the left-right shift remains.) Wavelength-dependent attenuation may introduce an additional term into the efficiency shown in figure 7, but no such detailed information is available. Since the fit described above is done using a generic model without direct use of wavelength-dependent details, this additional term has no effect on the analysis. We conservatively consider the extreme cases of all of the light having an attenuation length of 0.05 cm and all of it having an attenuation length of 0.125 cm. The difference in inferred I-value between these extremes is 0.7 eV. We take the midpoint as the nominal result, assume a uniform PDF over the range and claim an uncertainty of 0.2 eV.
@@ 5.4 Multiple Coulomb scattering
which is rapid to compute at the expense of accuracy. It predicts a longer projected range than EMZ for the same I-value, such that a fit to the data using Opt0 predictions results in an inferred I-value 5.5 eV lower than that inferred using EMZ. We do not use this rather different result except to note that experiments which rely on absolute particle ranges should certainly use EMZ and not the default Opt0.
@@ 6 Conclusions
Using our recommended value rather than the old values of 188 eV or 187 eV shifts energy reconstruction in liquid argon TPCs used for GeV-scale neutrinos by around 0.4%, with the exact shift depending on the analysis method. For instance, the dE/dx of a minimum-ionizing particle is reduced from 1.508 MeV g-1 cm2 to $(1.501\pm 0.002)$ MeV g-1 cm2. The inferred value of $\Delta m^{2}_{32}$ is directly proportional to reconstructed energy. As DUNE projects a sensitivity to this parameter of around 0.5% (depending on exposure) [32], is it important that a reliable value and uncertainty for the mean excitation energy of liquid argon be used.